We need help getting the answer to:
(6) (1 3/10) = _____
We think it is 7 4/5 but that is not one of the choices.
Can you help?

Answer by Nate(3500)

1 3/10 = 13/10
6(13/10) = 3(13/5) = 39/5 = 7 4/5